# FedGrAINS: Personalized SubGraph Federated Learning with AdaptIve Neighbor Sampling

### v1

@@ 1 Introduction
Graph Neural Networks (GNNs) operate over a single graph, with nodes and edges stored in a central server. For example, in social networks, each user, along with their connections form a large graph of all users and their links [2]. However, parties may not send their private graph datasets to a central server due to privacy concerns, raising the need to train GNN models over multiple distributed graph datasets. For instance, hospitals may maintain their patient interaction networks to track physical contacts or co-diagnoses of diseases. However, they cannot share these graphs with other entities due to privacy restrictions [52]. Here, the question is how to collaboratively train GNNs across clients’ distributed subgraphs without sharing the actual graph data. Federated Learning (FL) of GNNs aims to solve this problem so that each client individually trains a local GNN on their local data, while a central server aggregates locally updated GNN weights from multiple clients into a global model using federated optimization [19].
@@
On the other hand, dropout-based data pruning in FL[44, 21] and stochastic regularization in GNNs[30, 17] have shown to be successful in regularizing deeper models without any significant overhead by randomly dropping the network weights and/or data components. However, existing dropout-based data pruning approaches in FL are not suitable for graph ML models due to the invariance requirements of GNNs, forcing us to utilize dropout methods for graph-structured data [40, 11]. Furthermore, applying dropout techniques naively to deeper GNNs without accounting for the data structure can degrade the quality of node and structural information due to Laplacian smoothing, leading to over-smoothened representations that are uninformative in higher layers[31]. Moreover, using the same dropout parameters across different clients is also not an effective strategy due to heterogeneity over data, system, and the model, raising the need for coordination based on the users’ requirements.
To effectively personalize GNN models by respecting graph structures, we propose a novel data-adaptive, sampling-based regularization method for subgraph FL, named FedGrAINS by estimating and selecting essential nodes at each layer of GNNs every round for every client in a federated manner. To estimate the adaptive neighborhood distribution of each node in the datasets, we leverage Generative Flow Networks (GFlowNets), the state-of-the-art generative models tailored for structured discrete objects like graphs [7]. Here, for each node in the clients’ graph, we leverage GFlowNets to estimate the importance of the neighbors of a node so that we can prune the graph for each node adaptively in parallel to optimize the clients’ personalized task performances. We refer to this framework as Personalized SubGraph Federated Learning with AdaptIve Neighbor Sampling (FedGrAINS).
We extensively validate the performance of FedGrAINS across six popular graph datasets, evaluating overlapping and disjoint subgraph FL scenarios with varying numbers of clients. Our comprehensive experimental results show that FedGrAINS significantly outperforms relevant baselines. Our contributions are summarized as follows:
@@ 3.1 Adaptive Neighborhood Distribution Estimation with GFlowNets:
Using the modeling described above, in an $L$-layer GNN, we construct a sequence of $L$ adjacency matrices to reach a terminating state. We aim to perform task-aware sampling with the optimal GFlowNet sampling policy to minimize the expected classification loss :
@@
The GFlowNet of FedGrAINS is a neural network that learns the forward probabilities $P_{F}(s_{\ell+1}|s_{\ell})$, the probability of sampling an adjacency matrix $\bm{A}_{\ell+1}$ for the layer $\ell+1$ given the previous layers’ adjacency matrices $\bm{A}_{0},\dots,\bm{A}_{\ell}$ For each node $v_{i}$, where $i\in{1,\ldots,N}$, the GFlowNet models the probability that node $v_{i}$ in the neighborhood of the sampled nodes at layer $\ell$, $\mathcal{N}(\mathcal{S}^{\ell})$, will be included in $\mathcal{V}^{\ell+1}$, the set of nodes sampled for layer $\ell+1$. We represent the inclusion probability of each node in the sample using a Bernoulli distribution [11, 40]. In this formulation, $P_{F}(s_{l+1}|s_{l})$ becomes tractable, with each node represented as a Bernoulli-distributed random variable. Each node is associated with a logit indicating its likelihood of being included in the sampled set. Thus, if $p_{i}$ denotes the probability that the node $v_{i}$ in in the neighborhood of the sampled nodes at layer $\ell$ will be included in $\mathcal{V}^{\ell+1}$, then the forward probability transition function between layers $\ell$ and $\ell+1$ is $P_{F}(s_{\ell+1}|s_{\ell})=\prod_{v_{i}\in\mathcal{V}^{\ell+1}}p_{i}\prod_{v_{j}\in\mathcal{N}(\mathcal{S}^{\ell})\backslash\mathcal{V}^{\ell+1}}(1-p_{j})$, where the first product computes the probabilities of the nodes that are in the newly sampled set $\mathcal{V}^{\ell+1}$, and the second one computes the probabilities of nodes that are not within the newly sampled set $\mathcal{V}^{\ell+1}$.
@@
TB requires us to define the backward transition function $P_{B}(s_{\ell}|s_{\ell+1})$ [34]. In our setup, however, it is not required as the state representation, $s=(\bm{A}_{0},\dots,\bm{A}_{\ell})$, exactly saves the trajectory that was taken through the flow network to get to state $s$. Thus, the backward mapping is unique and deterministic. So, $P_{B}(s_{l}|s_{l+1})=1$.
@@ 3.1.1 Sampling and Off-Policy Training:
We want to sample $k$ important ones out of $N$ graph nodes without replacement at every message-passing step. However, our forward transition function $P_{F}$ contains $N$ independent Bernoulli distributions, and we are unlikely to sample exactly $k$ nodes from this distribution. Instead, we use the Gumbel-max trick [50, 29, 23], which selects a set of nodes $\mathcal{V}_{k}^{\ell}$ by perturbing the log probabilities randomly and taking the top-$k$ among those:
@@
where $G_{i}\sim\text{Gumbel}(0,1)$. The sampling process shown above differs from the sampling from the forward transition function $P_{F}$ sequentially one node at a time. Thus, we train the GFlowNet in an off-policy setting using the trajectory balance loss without any change [35]. In addition, the off-policy training is less expensive in computation and more robust to high-variance problems in distribution estimation than direct gradient-estimation methods such as REINFORCE[1]. Next, we explain how the adaptive importance sampling framework assists the (personalized) subgraph FL in Section 3.2.
@@ 4.2 Experimental Results
Tables 1 and 2 show node classification results in disjoint and overlapping subgraph scenarios, respectively. The primary motivation behind the proposed FedGrAINS is its applicability to existing subgraph FL algorithms. For a more fair comparison with Fed-PUB [4], we combine our data-adaptive regularization framework with their personalized aggregation and sparse masking. Other baselines do not include personalized aggregation. In addition, we did not test how the inclusion of our proposed framework affects the performance of FedSage+ as our proposed framework tries to filter unimportant nodes. In contrast, FedSage+ tries to increase the number of neighbors with fake neighbors, providing a more complicated setup for us to observe the effect of only learnable node filtering. Based on the experimental results, we make the following observations: First, regardless of the FL algorithm, classification performance decreases when the number of clients increases due to the increased number of partitions creating more missing links between the clients (See Appendix B.1.1 for the missing links counts for each dataset). The increasing number of missing links also leads to severe node-degree heterogeneity(see Appendix B.1.2 )within clients. This results in more challenging local training and collaboratively learning generalizable models with other clients.
@@
Finally, the inclusion of FedGrAINS as a personalization method consistently enhances the performance of all baselines. Specifically, the GFlowNet-based importance sampler improves the accuracy of all algorithms (except FedSage+) by at least 1%. Additionally, the advantages of FedGrAINS are most evident when comparing the results of FedAvg with our personalization method to those of FedPer [3], GCFL+ [49], and FedSage+ [52]. Moreover, the performance of both FedSage+ and FedAvg with FedGrAINS are very close for Cora, CiteSeer, and PubMed datasets. However, our personalized FL framework demonstrates significantly better performance for the Amazon-Computer, Amazon-Photo, and ogbn-arxiv datasets. We attribute this success to the size of the graph dataset, which resulted in a higher average degree than citation networks.
In comparison with FED-PUB [4], FED-PUB remains superior due to its use of personalized aggregation. To further assess the impact of FedGrAINS, we evaluate its performance when integrated with base algorithms that also employ personalized aggregation, such as FED-PUB, i.e., FED-PUB + FedGrAINS. Our results show that incorporating the importance sampler consistently and significantly enhances FED-PUB performance. Therefore, we can conclude that FedGrAINS can be seamlessly integrated into most subgraph FL algorithms without incurring additional communication overhead or compromising privacy.
@@
To understand the operation regime of our personalization method, we vary important hyperparameters for the GFlowNet-based sampler: the learning rate of the GFlowNet, $\beta_{GFN}$, and the reward scaling factor $\alpha$. We use the log uniform distribution to sample the aforementioned hyperparameters with the values from the following ranges respectively, $[1e^{-6},1e^{-2}]$ and $[1e^{2},1e^{6}]$. The setting is on the Cora dataset with the overlapping node scenario, where we set the number of local epochs to 1 and the number of clients to 10. The results are shown in Table 3. According to Table 3, the relationship between the performance and the reward scaling factor $\alpha$ is highly non-linear. In addition, the learning rate $\beta_{GFN}$ is not as effective as the scaling factor $\alpha$.

### v2

@@ 1 Introduction
Graph Neural Networks (GNNs) operate over a single graph, with nodes and edges stored in a central server. For example, in social networks, each user, along with their connections form a large graph of all users and their links [2]. However, parties may not send their private graph datasets to a central server due to privacy concerns, raising the need to train GNN models over multiple distributed graph datasets. For instance, hospitals may maintain their patient interaction networks to track physical contacts or co-diagnoses of diseases. However, they cannot share these graphs with other entities due to privacy restrictions [52]. Here, the question is how to collaboratively train GNNs across clients’ distributed subgraphs without sharing the actual graph data. Federated Learning (FL) of GNNs aims to solve this problem so that each client individually trains a local GNN on their local data, while a central server aggregates locally updated weights from multiple clients into a global model using federated optimization [19].
@@
On the other hand, dropout-based data pruning in FL[44, 21] and stochastic regularization in GNNs[30, 17] have shown to be successful in regularizing deeper models without any significant overhead by randomly dropping the network weights and/or data components. Existing dropout-based data pruning approaches in FL, however, are not suitable for graph ML models due to GNNs’ invariance requirements, which forces us to utilize dropout methods for graph-structured data [40, 11]. Further, applying dropout techniques naïvely to deeper GNNs without accounting for the data structure can degrade node and structural information quality due to Laplacian smoothing, leading to over-smoothened representations that are uninformative in higher layers[31]. Moreover, using the same dropout parameters across different clients is also not an effective strategy due to heterogeneity over data, system, and the model, raising the need for coordination based on the users’ requirements.
To effectively personalize GNN models by respecting graph structures, we propose a novel data-adaptive, sampling-based regularization method for subgraph FL, named FedGrAINS by estimating and selecting essential nodes at each layer of GNNs every round for every client in a federated manner. To estimate the adaptive neighborhood distribution of each node in the datasets, we leverage Generative Flow Networks (GFlowNets), the state-of-the-art generative models tailored for structured discrete objects like graphs [7]. Here, for each node in the clients’ graph, we leverage GFlowNets to estimate the importance of the neighbors of a node so that we can prune the graph for each node adaptively in parallel to optimize the clients’ personalized task performances. We refer to this framework as Personalized SubGraph Federated Learning with AdaptIve Neighbor Sampling (FedGrAINS). We extensively validate the performance of FedGrAINS across six popular graph datasets, evaluating overlapping and disjoint subgraph FL scenarios with varying numbers of clients. Our comprehensive experimental results show that FedGrAINS significantly outperforms relevant baselines. Our contributions are summarized as follows:
@@ 3.1 Adaptive Neighborhood Distribution Estimation with GFlowNets:
Using the modeling described above, in a $L$-layer GNN, we construct a sequence of $L$ adjacency matrices to reach a terminating state. We aim to perform task-aware sampling with the optimal GFlowNet sampling policy to minimize the expected classification loss :
@@
The GFlowNet of FedGrAINS is a neural network that learns the forward probabilities $P_{F}(s_{\ell+1}|s_{\ell})$, the probability of sampling an adjacency matrix $\bm{A}_{\ell+1}$ for the layer $\ell+1$ given the previous layers’ adjacency matrices $\bm{A}_{0},\dots,\bm{A}_{\ell}$ For each node $v_{i}$, where $i\in{1,\ldots,N}$, the GFlowNet models the probability that node $v_{i}$ in the neighborhood of the sampled nodes at layer $\ell$, $\mathcal{N}(\mathcal{S}^{\ell})$, will be included in $\mathcal{V}^{\ell+1}$, the set of nodes sampled for layer $\ell+1$. We represent the inclusion probability of each node in the sample using a Bernoulli distribution [11, 40]. In this formulation, $P_{F}(s_{l+1}|s_{l})$ is tractable thanks to Bernoulli parameterization for each node associated with a logit indicating its likelihood of being included in the sampled set. Thus, if $p_{i}$ denotes the probability that node $v_{i}$ in the neighborhood of the sampled nodes at layer $\ell$ will be included in $\mathcal{V}^{\ell+1}$, then the forward probability transition function between layers $\ell$ and $\ell+1$ is $P_{F}(s_{\ell+1}|s_{\ell})=\prod_{v_{i}\in\mathcal{V}^{\ell+1}}p_{i}\prod_{v_{j}\in\mathcal{N}(\mathcal{S}^{\ell})\backslash\mathcal{V}^{\ell+1}}(1-p_{j})$, where the first product computes the probabilities of the nodes that are in the newly sampled set $\mathcal{V}^{\ell+1}$, and the second one computes the probabilities of nodes that are not within the newly sampled set $\mathcal{V}^{\ell+1}$.
@@
TB requires us to define the backward transition function $P_{B}(s_{\ell}|s_{\ell+1})$ [34]. However, in our setup, it is not required as the state representation, $s=(\bm{A}_{0},\dots,\bm{A}_{\ell})$, exactly saves the trajectory that was taken through the flow network to get to the state $s$. Thus, the backward mapping is unique and deterministic. So, $P_{B}(s_{l}|s_{l+1})=1$.
@@ 3.1.1 Sampling and Off-Policy Training:
We want to sample $k$ important ones from $N$ graph nodes without replacement at every message passing step. However, our forward transition function $P_{F}$ contains independent Bernoulli distributions $N$, and it is unlikely that we will sample exactly $k$ nodes of this distribution. Instead, we use the Gumbel-max trick [50, 29, 23], which selects a set of nodes $\mathcal{V}_{k}^{\ell}$ by perturbing the log probabilities randomly and taking the top-$k$ among those:
@@
where $G_{i}\sim\text{Gumbel}(0,1)$. The sampling process shown above differs from the sequential sampling of the forward transition function $P_{F}$ one node at a time. Thus, we train the GFlowNet in an off-policy setting using the trajectory balance loss without any change [35]. In addition, the off-policy training is less expensive in computation and more robust to high-variance problems in distribution estimation than direct gradient-estimation methods such as REINFORCE[1]. Next, we explain how the adaptive importance sampling framework assists the (personalized) subgraph FL in Section 3.2.
@@ 4.2 Experimental Results
Tables 1 and 2 show node classification results in disjoint and overlapping subgraph scenarios, respectively. The primary motivation behind the proposed FedGrAINS is its applicability to existing subgraph FL algorithms. For a more fair comparison with Fed-PUB [4], we combine our data-adaptive regularization framework with their personalized aggregation and sparse masking. Other baselines do not include personalized aggregation. In addition, we did not test how the inclusion of our proposed framework affects the performance of FedSage+ as our proposed framework tries to filter unimportant nodes. In contrast, FedSage+ tries to increase the number of neighbors with fake neighbors, providing a more complicated setup for us to observe the effect of only learnable node filtering. Based on the experimental results, we make the following observations: First, regardless of the FL algorithm, classification performance decreases when the number of clients increases due to the increased number of partitions creating more missing links between the clients (See Appendix B.1.1 for the missing links counts for each dataset). The increasing number of missing links also leads to severe node-degree heterogeneity(see Appendix B.1.2 )within clients. This results in more challenging local training and collaborative learning of generalizable models with other clients.
@@
Finally, the inclusion of FedGrAINS as a personalization method consistently improves the performance of all baselines. Specifically, the GFlowNet-based importance sampler improves the accuracy of all algorithms (except FedSage+) by at least 1%. Additionally, the advantages of FedGrAINS are most evident when comparing the results of FedAvg with our personalization method to those of FedPer [3], GCFL+ [49], and FedSage+ [52]. Moreover, the performance of both FedSage+ and FedAvg with FedGrAINS are very close for the Cora, CiteSeer, and PubMed datasets. However, our personalized FL framework demonstrates significantly better performance for Amazon-Computer, Amazon-Photo, and ogbn-arxiv datasets. We attribute this success to the size of the graph dataset, which resulted in a higher average degree than citation networks.
Compared to FED-PUB [4], FED-PUB remains superior due to its use of personalized aggregation. To further assess the impact of FedGrAINS, we evaluate its performance when integrated with base algorithms that also employ personalized aggregation, such as FED-PUB, i.e., FED-PUB + FedGrAINS. Our results show that incorporating the importance sampler consistently and significantly enhances FED-PUB performance. Therefore, we can conclude that FedGrAINS can be seamlessly integrated into most subgraph FL algorithms without incurring additional communication overhead or compromising privacy.
@@
To understand the operation regime of our personalization method, we vary important hyperparameters for the GFlowNet-based sampler: the learning rate of the GFlowNet, $\beta_{GFN}$, and the reward scaling factor $\alpha$. We use the log-uniform distribution to sample the aforementioned hyperparameters with the values from the following ranges, respectively, $[1e^{-6},1e^{-2}]$ and $[1e^{2},1e^{6}]$. The setting is on the Cora dataset with the overlapping node scenario, where we set the number of local epochs to 1 and the number of clients to 10. The results are shown in Table 3. According to Table 3, the relationship between performance and reward scaling factor $\alpha$ is highly non-linear. In addition, the learning rate $\beta_{GFN}$ is not as effective as the scaling factor $\alpha$.